OpenStudy (anonymous):

Write an equivalent expression for that does not contain powers of trigonometric functions greater than 1 of : sin^4 x Help please

OpenStudy (anonymous):

It's sin to the 4th power btw not sin to the 4x power the 4 and x are separate

OpenStudy (larseighner):

\[ \Large \sin^2\theta = {{1 - \cos 2\theta} \over 2} \]